WHAT WILL BE THE CONCENTRATION OF CHLORIDE ION IF 200 ml OF  0.2 M CaCl2 IS MIXED WITH 300 ml OF 0.1 M NaCl SOLUTION ?

Grade:12

1 Answers

Arnab
6 years ago
I think you know this formula, M1V1+M2V2=(V1+V2)MUsing this formula, 200*0.2+300*0.1=(300+200)*Mor, M=0.14